My math problem is to find the discriminant of y=x^2-3x-5.
Could you please help me??

Answer by stanbon(75887)

find the discriminant of y=x^2-3x-5
a = 1, b = -3, c = -5
-------------
b^2 - 4ac = 9 -4*1*-5 = 9 + 20 = 29
===================================
Cheers,
Stan H.